The force required to accelerate a car weighing 2000 kg from rest to a speed of 30 m / s during 10s is: 

 A) 3x10⁴ N B) 6x10³ N C) 1200 N   D) 500 N


1
Expert's answer
2021-03-14T19:19:15-0400
"a=\dfrac{v-v_0}{t}=\dfrac{30\ \dfrac{m}{s}-0}{10\ s}=3\ \dfrac{m}{s^2},""F=ma=2000\ kg\cdot3\ \dfrac{m}{s^2}=6\cdot10^{3}\ N."

Answer:

B) 6x10³ N.